# Generation and Preservation of Large Entangled States on Physical Quantum Devices

### v1

@@ IV Efficient Bipartite Entanglement Characterisation in Whole-Device Graph States
Graph states are a convenient choice for studying large-scale entanglement as they are simple to prepare and comparitively noise robust [51]. Controlled-phase operations that do not overlap vertices can be applied in parallel, allowing any graph state to be prepared by a linear-size constant-depth circuit [57]. Concretely, a bounded degree graph can be prepared with a two-qubit gate depth equal to the maximum degree between its vertices [58]. An example least-depth graph state preparation circuit on a seven-qubit layout is shown in Fig. 10. The qubit layout is represented in Fig. 10a where nodes represent qubits and edges display possible CNOT operations. Since the graph has a maximum node degree of three, its graph state preparation circuit has a two-qubit gate depth of three.
[…]
To characterise bipartite entanglement across an entire device, we prepare a native-graph state containing every edge. We then perform full quantum state tomography on every locally entangled bipartition corresponding to each edge on the device. In detail, graph states have the property that projecting all but two qubits in an entangled cluster leaves the pair in a maximally entangled Bell state [59]. The extent of two-qubit entanglement can then be quantified by measuring the negativity [60]. For a quantum state represented by the density matrix , the negativity between subsystems and is calculated as
where are the eigenvalues of , where is the partial transpose of with respect to subsystem . A maximally entangled Bell state has a negativity of , whereas a fully separable state has a negativity of . Although there are many other entanglement measures, such as entropy of entanglement and concurrence, negativity is chosen because it is easy to compute and is related to the teleportation distance [60]. Specifically, the minimum teleportation distance achievable with state , acting on , is bound by
@@
Bipartite Entanglement Characterisation Protocol
We develop a protocol to efficiently characterise bipartite entanglement on quantum computers, inspired by experiments from Mooney et al. [22, 23]. Development of the scheme is driven by three main design principles: to devise an entanglement characterisation protocol that is highly automated, scalable, and architecture-independent. We implement the protocol in Python, utilising the Qiskit API to interface with IBM superconducting quantum devices. However, the techniques and procedures in the program are generally applicable. The protocol can be divided into five components:
@@
IV.0.1 Native-Graph State Preparation
@@
IV.0.2 Parallel Quantum State Tomography
In previous graph state experiments [22, 23], full quantum state tomography (QST) is performed on one qubit pair (and its neighbours) at a time. Using such method, the number of circuits required to fully characterise a device increases linearly with the number of edges. Our procedure improves upon this by executing QST in parallel. In detail, we perform simultaneous basis measurements on non-overlapping sets of qubits, where each set defines a target qubit pair and its neighbours. By grouping sets into batches and performing parallel QST for batches at a time, we can fully characterise bipartite entanglement on any size device (provided invariable qubit topology) with a constant number of circuits.
Using a specialised scheduling scheme for the heavy hex hardware layout, the number of required batches can be reduced to only four when allowing Bell pairs to share neighbours. This can be achieved by using an alternate tiling of two unit cells (able to be rotated 180°), where each unit cell is composed of six edges and each edge is assigned to one of four batches. However, similar to the case of graph state embedding, we instead opt for a topology-agnostic algorithm that performs parallel QST for as many non-overlapping Bell pairs as it can fit into a single batch. In our implementation, we prevent target pairs from sharing neighbours for practical convenience. Table 2 lists the number of batches required to characterise IBM Quantum devices up to 433 qubits in size.
@@ IV.0.2 Parallel Quantum State Tomography
Bipartite Entanglement on IBM Quantum Devices
We characterise bipartite entanglement on all IBM Quantum devices accessible by the University of Melbourne IBM Quantum Hub. At the time of experiment, these include four 5-qubit systems, five 7-qubit systems, one 16-qubit system, eight 27-qubit systems, three 127-qubit systems, and one 433-qubit system, totaling to 22 systems. Both unmitigated and mitigated results using QREM are shown. It is important to present unmitigated results because not all protocols involving graph states can incorporate readout error mitigation. Notably, quantum teleportation schemes which use mid-circuit measurements are incompatible with QREM.
@@
We summarise all graph state experiment results in Table 3. In addition to tabulating device size, quantum volume, and mean device negativity, we display the sizes of the largest connected entanglement graphs with edges above a certain negativity threshold. In detail, columns with label represent the size of the largest connected graph where edges exist only between qubits pairs whose measured negativity is at least of the maximum value. This metric allows us to simultaneously probe the scale and quality of clusters of entanglement. We observe a few general trends. Firstly, the standard deviation in negativities for each device typically decreases once we apply QREM. Similarly, the variance between mean device negativities also diminishes. This may be attributed to the bound on the maximum negativity and variance in mean readout error rates, which range from 1.1% on ibm_lagos to 5.2% on ibmq_quito. Secondly, for devices of similar size, quantum volume is not a good predictor of mean device negativity. For instance, ibm_geneva, which has a quantum volume of 32, has a mean device negativity of (QREM) compared to ibm_kolkata’s , which has a quantum volume of 128. This may be ascribed to a couple of factors, the first being that quantum volume is defined over a subset of qubits, instead of the whole device, and the second being that our entanglement protocol utilises primarily low-depth circuits.
@@
seattle | 433 | - | 330 | 37 | 11 | ✓ (active qubits) |
Generating 414-Qubit Graph States
Using the same protocol, we characterise bipartite entanglement on a larger 414-qubit graph state prepared on the 433-qubit device , where at the time of experiment, 19 of the 433 device qubits were inoperable. Figure 17 displays the average negativity versus nearest-neighbour qubit pairs. The average qubit pair negativity is found to be 0.115 without QREM, and 0.340 with QREM. We report proportionally higher readout error rates compared to previous devices. The coupling map is displayed in Fig. 18. After mitigating for readout errors, all bipartitions not involving inactive qubits had measured negativities above 0.
@@ V Preserving Whole-Device Graph States Via Dynamical Decoupling
The methodology for testing dynamical decoupling on whole-device graph states parallels the GHZ case. A whole-device graph state is prepared as previously outlined. A variable delay period is inserted between state preparation and whole-device quantum state tomography. We compare negativity decays for circuits without mitigation (free decay), and circuits mitigated with PDD and Hahn echo. The general circuit diagram is shown in Fig. 19 and the PDD circuit in Fig. 6c.
[…]
Whole-Device Graphstate Negativity Decays on a 127-Qubit Device
@@
While we focus on mean improvements in entanglement lifetimes for native graph states, recent results have shown that implementing a precisely timed dynamical decoupling sequence can more effectively cancel the coherent ZZ errors in a 12-qubit ring graph state [61]. A similar approach of tailoring the dynamical decoupling sequence for native graph states may also be beneficial, although can be considerably more complex depending on the scale and connectivity of the underlying graph. For the purposes of improving the mean entanglement lifetime of a large-scale graph state, we show that even a simple PDD sequence incurs significant benefits. Additionally, we note that an adaption of our procedure may be potentially useful in detecting and characterising the coherent noise effects.
@@ VI Discussion
For graph states, we developed a bipartite entanglement characterisation protocol that constructs entanglement graphs depicting bipartite entanglement in IBM Quantum devices using as low as a constant 36 circuits. We used the protocol to verify and quantify whole-device bipartite entanglement over 20 different IBM Quantum systems, including three 127-qubit systems. We further showed entanglement across 414 qubits in a 433-qubit device. We then tested dynamical decoupling for preserving qubit pair negativities in a native-graph state prepared on the 127-qubit ibmbrisbane device. We observed coherent noise signals consistent with residual ZZ interactions, which were partially suppressed after application of PDD. PDD led to an overall improvement in mean device bipartite entanglement lifetimes. We also note the potential utility of a running a similar procedure to detect and characterise the coherent noise signals.

### v2

@@ IV.1 Graph States
Graph states are a convenient choice for studying large-scale entanglement as they are simple to prepare and comparitively noise robust [52]. Controlled-phase operations that do not overlap vertices can be applied in parallel, allowing any graph state to be prepared by a linear-size constant-depth circuit [58]. Concretely, a bounded degree graph can be prepared with a two-qubit gate depth equal to the maximum degree between its vertices [59]. An example least-depth graph state preparation circuit on a seven-qubit layout is shown in Fig. 10. The qubit layout is represented in Fig. 10(a) where nodes represent qubits and edges display possible CNOT operations. Since the graph has a maximum node degree of three, its graph state preparation circuit has a two-qubit gate depth of three.
To characterize bipartite entanglement across an entire device, we prepare a native-graph state containing every edge. We then perform full quantum state tomography on every locally entangled bipartition corresponding to each edge on the device. In detail, graph states have the property that projecting all but two qubits in an entangled cluster leaves the pair in a maximally entangled Bell state [60]. The extent of two-qubit entanglement can then be quantified by measuring the negativity [61]. For a quantum state represented by the density matrix , the negativity between subsystems and is calculated as
where are the eigenvalues of , where is the partial transpose of with respect to subsystem . A maximally entangled Bell state has a negativity of , whereas a fully separable state has a negativity of . Although there are many other entanglement measures, including more complex multipartite entanglement witnesses, negativity is an entanglement monotone that is simple to compute. A non-zero measurement for the negativity on a 2-qubit state is a necessary and sufficient condition for entanglement. This makes it a great choice for bipartite entanglement on graph states, where edges can be reduced to Bell states. Additionally, negativity is related to the minimum teleportation distance achievable with state acting on :
@@
IV.2 Bipartite Entanglement Characterization Protocol
We develop a protocol to efficiently characterize bipartite entanglement on quantum computers, inspired by experiments from Mooney et al. [22, 23]. Development of the scheme is driven by three main design principles: to devise an entanglement characterization protocol that is highly automated, scalable, and architecture-independent. We implement the protocol in Python, utilizing the Qiskit API to interface with IBM superconducting quantum devices. However, the techniques and procedures in the program are generally applicable. The protocol can be divided into five components:
@@ IV.2 Bipartite Entanglement Characterization Protocol
IV.2.1 Native-Graph State Preparation
@@
IV.2.2 Parallel Quantum State Tomography
In previous graph state experiments [22, 23], full quantum state tomography (QST) is performed on one qubit pair (and its neighbours) at a time. Using such method, the number of circuits required to fully characterize a device increases linearly with the number of edges. Our procedure improves upon this by executing QST in parallel. In detail, we perform simultaneous basis measurements on non-overlapping sets of qubits, where each set defines a target qubit pair and its neighbours. By grouping sets into batches and performing parallel QST for batches at a time, we can fully characterize bipartite entanglement on any size device (provided invariable qubit topology) with a constant number of circuits.
Using a specialized scheduling scheme for the heavy hex hardware layout, the number of required batches can be reduced to only four when allowing Bell pairs to share neighbours. This can be achieved by using an alternate tiling of two unit cells (able to be rotated 180°), where each unit cell is composed of six edges and each edge is assigned to one of four batches. However, similar to the case of graph state embedding, we instead opt for a topology-agnostic algorithm that performs parallel QST for as many non-overlapping Bell pairs as it can fit into a single batch. In our implementation, we prevent target pairs from sharing neighbours for practical convenience. Table 3 lists the number of batches required to characterize IBM Quantum devices up to 433 qubits in size.
@@ IV.2.2 Parallel Quantum State Tomography
IV.3 Bipartite Entanglement on IBM Quantum Devices
We characterize bipartite entanglement on all IBM Quantum devices accessible by the University of Melbourne IBM Quantum Hub. At the time of experiment, these include four 5-qubit systems, five 7-qubit systems, one 16-qubit system, eight 27-qubit systems, three 127-qubit systems, and one 433-qubit system, totaling to 22 systems. Both unmitigated and mitigated results using QREM are shown. It is important to present unmitigated results because not all protocols involving graph states can incorporate readout error mitigation. Notably, quantum teleportation schemes which use mid-circuit measurements are incompatible with QREM.
@@ IV.3 Bipartite Entanglement on IBM Quantum Devices
We summarize all graph state experiment results in Table 4. In addition to tabulating device size, quantum volume, and mean device negativity, we display the sizes of the largest connected entanglement graphs with edges above a certain negativity threshold. In detail, columns with label represent the size of the largest connected graph where edges exist only between qubits pairs whose measured negativity is at least of the maximum value. This metric allows us to simultaneously probe the scale and quality of clusters of entanglement. We observe a few general trends. Firstly, the standard deviation in negativities for each device typically decreases once we apply QREM. Similarly, the variance between mean device negativities also diminishes. This may be attributed to the bound on the maximum negativity and variance in mean readout error rates, which range from 1.1% on ibm_lagos to 5.2% on ibmq_quito. Secondly, for devices of similar size, quantum volume is not a good predictor of mean device negativity. For instance, ibm_geneva, which has a quantum volume of 32, has a mean device negativity of (QREM) compared to ibm_kolkata’s , which has a quantum volume of 128. This may be ascribed to a couple of factors, the first being that quantum volume is defined over a subset of qubits, instead of the whole device, and the second being that our entanglement protocol utilizes primarily low-depth circuits.
@@
| seattle | 433 | - | 330 | 37 | 11 | ✓ (active qubits) | |
IV.4 Generating 414-Qubit Graph States
Using the same protocol, we characterize bipartite entanglement on a larger 414-qubit graph state prepared on the 433-qubit device , where at the time of experiment, 19 of the 433 device qubits were inoperable. Figure 17 displays the average negativity versus nearest-neighbour qubit pairs. The average qubit pair negativity is found to be 0.115 without QREM, and 0.340 with QREM. We report proportionally higher readout error rates compared to previous devices. The coupling map is displayed in Fig. 18. After mitigating for readout errors, all bipartitions not involving inactive qubits had measured negativities above 0.
@@ V Preserving Whole-Device Graph States Via Dynamical Decoupling
The methodology for testing dynamical decoupling on whole-device graph states parallels the GHZ case. A whole-device graph state is prepared as previously outlined. A variable delay period is inserted between state preparation and whole-device quantum state tomography. We compare negativity decays for circuits without mitigation (free decay), and circuits mitigated with PDD and Hahn echo. The general circuit diagram is shown in Fig. 19 and the PDD circuit in Fig. 6(b).
V.1 Whole-Device Graphstate Negativity Decays on a 127-Qubit Device
@@
While we focus on mean improvements in entanglement lifetimes for native graph states, recent results have shown that implementing a precisely timed dynamical decoupling sequence can more effectively cancel the coherent ZZ errors in a 12-qubit ring graph state [62]. A similar approach of tailoring the dynamical decoupling sequence for native graph states may also be beneficial, although can be considerably more complex depending on the scale and connectivity of the underlying graph. For the purposes of improving the mean entanglement lifetime of a large-scale graph state, we show that even a simple PDD sequence incurs significant benefits. Additionally, we note that an adaption of our procedure may be potentially useful in detecting and characterizing the coherent noise effects.
@@ VI Discussion
For graph states, we developed a bipartite entanglement characterization protocol that constructs entanglement graphs depicting bipartite entanglement in IBM Quantum devices using as low as a constant 36 circuits. We used the protocol to verify and quantify whole-device bipartite entanglement over 20 different IBM Quantum systems, including three 127-qubit systems. We further showed entanglement across 414 qubits in a 433-qubit device. We then tested dynamical decoupling for preserving qubit pair negativities in a native-graph state prepared on the 127-qubit ibmbrisbane device. We observed coherent noise signals consistent with residual ZZ interactions, which were partially suppressed after application of PDD. PDD led to an overall improvement in mean device bipartite entanglement lifetimes. We also note the potential utility of a running a similar procedure to detect and characterize the coherent noise signals.
@@ Acknowledgements.
- Cross et al. [2019] A. W. Cross, L. S. Bishop, S. Sheldon, P. D. Nation, and J. M. Gambetta, Validating quantum computers using randomized model circuits, Physical Review A 100, 10.1103/PhysRevA.100.032328 (2019).